# A Hierarchy of Spectral Gap Certificates for Frustration-Free Spin Systems

### 06 Nov 2024 v1

@@ I Introduction
Proving the existence of a spectral gap is a highly non-trivial mathematical challenge. Even for reasonable local Hamiltonians the task can be undecidable [11, 12, 13]. Several techniques to lower bound the spectral gap have been developed for the better-behaved class of frustration-free Hamiltonians, i.e. systems in which the ground state minimizes each local term individually. The most notable lower bounds are based on the “martingale” method [14] and finite size criteria [15, 16]. More recently, methods inspired by the numerical bootstrap approach to conformal field theory have been proposed to determine spectral properties of simple quantum Hamiltonians as well as quantum many body systems [17, 18, 19].
@@
To put our results into context, we find it worthwhile to pause and briefly discuss the basic idea behind our method—and how it generalizes existing gap certification methods—on a slightly more technical level. The well known idea underlying all gap certification approaches is the fact that a frustration-free Hamiltonian $H$ has a gap greater or equal to $\delta$ if the operator $H^{2}-\delta H$ is positive semidefinite (i.e. all its eigenvalues are non negative). We notice that all of the above mentioned gap certification methods can be regarded as finding different ways to decompose $H^{2}-\delta H$ as a sum of positive terms. Inspecting the different known methods, we notice that in order to attain better lower bounds it seems to be necessary to allow for more freedom in choosing the form of the terms in the decomposition. A good example of this is Gosset and Mozgunov’s [20] improvement over Knabe’s bound [15] which was obtained by introducing additional free parameters in what is otherwise a very similar proof. This observation leads us to formulate the gap estimation problem as an optimization problem where $\delta$ is maximized under the constraint that $H^{2}-\delta H$ can be written as a sum of positive terms. A first version of this idea has been proposed in Ref. [37]. To exhaust the search space in a systematic way we organize it as hierarchy of optimization problems: at each level of the hierarchy we optimize over all $n$-local positive decompositions of $H^{2}-\delta H$.
The rest of the paper is organized as follows: In Section II, we discuss the main technical ingredients of our method. In Section III, we show how previous finite-size criteria arise as special cases of the semidefinite program introduced in Section II. In Section IV, we present numerical results comparing lower bounds obtained using our method against those obtained with existing finite-size criteria.
II The method
@@
Without loss of generality, we will shift the ground state energy to be 00, which, combined with frustration-freeness, implies both $H\succeq 0$ and $h_{i}\succeq 0$, where $\succeq$ is used to denote that the matrix is positive semidefinite.
The spectral gap is then the lowest non-zero eigenvalue of $H$ and we denote it by $\Delta_{N}^{\circ}$, where the $\circ$ superscript symbolizes periodic boundary conditions. Of particular relevance is the behavior of $\Delta_{N}^{\circ}$ as the system size increases. A model is defined to be gapped if the spectral gap does not close in the thermodynamic limit, that is $\Delta_{N\to\infty}^{\circ}=\Delta>0$.
@@
The above condition enforces that $H$ has no eigenvalues in the interval $(0,\delta)$. Let us stress that all the most commonly-used methods to prove the gap of a system rely on establishing condition (2) for a given value of $\delta$ [15, 23, 14, 21, 36, 34, 35].
The condition in eq. 2 allows us to formulate the gap estimation problem as an optimization problem, which truns out to be a semidefinite program (SDP).
$\Delta_{N}^{\circ}$ $=\max\delta\,\,$
@@
under the condition that the ground state energy of $H$ is zero. The optimal value of this SDP is exactly the energy of the first excited state of $H$. Additionally, at the optimal solution, both the ground state and the first excited state space of $H$ lie in the ground state space of $H^{2}-\Delta H$.
In the general case, solving this SDP is as hard as exactly diagonalizing the Hamiltonian, since this problem imposes a positivity constraint on a matrix of size $(d^{N}\times d^{N})$. As $N$ increases, very quickly it becomes computationally intractable to solve. In the following, we show how to relax this problem. The resulting relaxation is a local, $n$-body, problem the solution to which is a lower bound on the gap of $N$ site periodic system for all $N\geq 2n$, and therefore holds for the infinite system as well.
II.1 Simplifications of the general SDP
@@
Geometric Locality condition. The first problem we need to address is that $H^{2}$ is not local: it consists of products of interactions terms $h_{i},h_{j}$ arbitrarily far apart. We therefore introduce a truncated operator $\left[H^{2}\right]_{n}$ which is $n$-local, by removing terms that act on spins farther than $n$ sites apart, namely
$\begin{split}\left[H^{2}\right]_{n}&=\sum_{|i-j|<n}h_{i}h_{j}\\ &=\sum_{i}\left(h_{i}^{2}+\sum_{i<j<i+n-k}\{h_{i},h_{j}\}\right).\end{split}$ (4)
If $n$ is chosen suitably (i.e. $n\geq 2k+1$), terms $h_{i}$ and $h_{j}$ with $j\geq i+n$ act on disjoint sets of spins and thus every discarded term is positive semidefinite (as a tensor product of positive terms). It follows that, if we define the following operator,
@@
where the the $i$ subscript indicates the position of the first spin on which $q_{n}$ acts (the $n$ subscript simply denotes the $n$-locality, i.e. that $q_{n}$ is supported on $n$ neighboring sites). If for some value of $\delta$ $Q_{n}(\delta)$ is positive semidefinite, we can hope to find such a positive generating term because $Q_{n}(\delta)$ is itself $n$-local and translational invariant.
@@
We will now proceed to show how the condition (6) can be expressed in terms of local operators, and thus arrive at a finite size relaxation of section II that is numerically tractable for any finite support $n$.
II.2 Primal SDP
After substituting the operator $H^{2}$ in the exact SDP section II with $[H^{2}]_{n}$, and imposing the condition that $Q_{n}(\delta)$ has a positive operator decomposition we arrive at the following problem
@@ II.2 Primal SDP
where $Y_{n-1}$ can be any operator supported on $n-1$ sites. Condition (10) is chosen such that, when summed over $i$, the terms involving $Y_{n-1}$ cancel telescopically. It follows that the constraint (7) corresponds to looking for a suitable $Y_{n-1}$ that makes the local generating term $q_{n,i}$ positive semidefinite. In this way, we arrive at the final form of our SDP, which reads
@@
We will refer to the problem we arrived at in eq. 11 as the locally translationally invariant (LTI) gap SDP. It is a semidefinite program which involves a positivity constraint on an $n$-body (i.e. $d^{n}\times d^{n}$ dimensional) matrix.
@@
II.3 Dual SDP
Any semidefinite program which is defined as a maximization has a dual formulation in terms of a minimization problem [38]. Importantly, the optimum of the dual serves as an upper bound on the primal optimal. When the two optima coincide, the problem exhibits so-called strong duality and one can equivalently solve either one of the problems. In this subsection we analyse the dual problems related to the spectral gap SDPs introduced in the previous subsections and show how they can provide quantitative information about the first excited state.
The dual program of the exact spectral gap SDP of section II is
@@ II.3 Dual SDP
This problem has strong duality, because the choice $\rho=\ket{\psi_{1}}\bra{\psi_{1}}/\Delta$, where |$\ket{\psi_{1}}$ denotes the first excited state of $H$; is feasible and attains the solution $\Delta$, which is the optimal solution of the primal. An alternate way to show that the optimal solution for the dual is directly related to the first excited state is highlighted in Appendix B.
This formulation of the gap problem provides a physical interpretation to the truncation step in eq. 4, where we disregarded far away terms $\{h_{i},h_{j}\}$ in $H^{2}$. As we see in eq. 15 the exact gap is given by the expectation value of $H^{2}$ in the first excited state. If that state exhibits fast decaying correlations then the contribution of the far away terms we neglected should not noticeably change the result.
Analogously to eq. 15, the dual of the TI-SDP problem has the form
@@
where the optimization variable $\rho_{n}$ corresponds to the reduced marginals on $n$ parties. The constraint
@@
is called local translation invariance (LTI) condition and arises as a dual to the constraint eq. 12. Moreover due to strong duality, the following relation is satisfied between the primal and dual optimal variables,
@@
This relation which is known as complementarity slackness has several useful consequences. Among others, it provides us with the gradient of the solution to the SDP with respect to parameters in the problem data, without incurring further computational cost [39]. In our case the data is the Hamiltonian interaction term. We can then use the gradient to optimize the gap over all parent Hamiltonians of a given ground state. We elaborate on this in Appendix D.
@@
III Connection to previous finite-size criteria
In this section, we compare our method to finite-size criteria-based techniques, which are the prevalent methods used to lower bound spectral gaps. For consistency with the formulation of the existing methods[15, 20], we restrict the discussion to nearest-neighbor projector Hamiltonians
@@ III Connection to previous finite-size criteria
Importantly, if the local term $h$ of a given frustration-free Hamiltonian is not a projector, it can always be modified into a projector Hamiltonian by adjusting the eigenvalues of the subspace orthogonal to the ground space of $h$. It follows from a simple argument that proving the gap of this second Hamiltonian implies a gap for the first as well [20](c.f. Section IV). We will now prove that the set of feasible local terms in the LTI-SDP eq. 11 includes the ones obtained from known finite size criteria. Thus, the value of LTI SDP achieves a tighter lower bound on the spectral gap.
III.1 Finite-size criteria
@@
(Method 1) Based on finite system gaps. First formulated by Knabe [15], these criteria provide a way to relate the spectral gap of a finite open-boundary system, to the spectral gap in the thermodynamic limit. The open boundary Hamiltonian on a finite system of size $n$ is,
$h_{n,i}=\sum\limits_{j=0}^{n-1}h_{i+j},$ (20)
and $i$ denotes the site index where the support of the open boundary Hamiltonian begins. If $h_{n,i}$ has a spectral gap of $\epsilon_{n}$, then the lower bound to the gap in the thermodynamic limit is given as
$\Delta(n)=f(n)\left(\epsilon_{n}-g(n)\right).$ (21)
If $\Delta_{m}^{\circ}$ is the gap of periodic system on $m$ sites, then the Knabe bound [15] is
$\Delta_{m}^{\circ}\geq\Delta_{\mathrm{Knabe}}(n)=\frac{n-1}{n-2}\left(\epsilon_{n}-\frac{1}{n-1}\right)$ (22)
@@
$\Delta_{m}^{\circ}\geq\Delta_{\mathrm{GM}}(n)=\frac{5}{6}\frac{n^{2}+n}{n^{2}-4}\left(\epsilon_{n}-\frac{6}{n^{2}+n}\right)$ (23)
for $m>2n>4$. Comparing the bounds with eq. 21, the functions $f(n)$ and $g(n)$ satisfy the following properties
$\lim_{n\to\infty}f(n)\leq 1,\quad\lim_{n\to\infty}g(n)=0.$ (24)
@@
Moreover, for the case of gapless models, the function $g(n)$ provides an upper bound on spectral gap scaling with $n$ for the open boundary Hamiltonian $h_{n,i}$. The Gosset–Mozgunov bound improved this upper bound to $\mathcal{O}(n^{-2})$ from $\mathcal{O}(n^{-1})$ concluded from the Knabe bound. In a recent article, this $\mathcal{O}(n^{-2})$ scaling was generalized to arbitrary graphs and lattice dimensions [41].
@@
This method was introduced by Nachtergaele, Fannes, and Werner in Ref. [14, 16], and is also known as the ”martingale” criterion. Instead of using the gap of a finite size system, the presence of a gap is established by checking if a relation in terms of the ground space projectors is satisfied. The bound is obtained by first constructing a new Hamiltonian $\tilde{H}=\sum_{i}\tilde{h}_{\mathcal{I},\mathcal{I}+1}$ obtained by blocking $l$ sites into a single site, i.e.
$\{i,...,i+l-1\}$ $\to\mathcal{I}$ (26)
$\tilde{h}_{\mathcal{I},\mathcal{I}+1}=\sum\limits_{x=0}^{l+1}h_{i(l+1)-l+x}$ $\Rightarrow H\geq\frac{\tilde{H}}{2}.$ (27)
W.l.o.g, the blocked Hamiltonian $\tilde{h}$ can be converted into the corresponding projector Hamiltonian $K=\sum_{i}P_{\mathcal{I},\mathcal{I}+1}$, which satisfies the relation,
$\tilde{h}_{\mathcal{I},\mathcal{I}+1}\geq\epsilon_{2l}P_{\mathcal{I},\mathcal{I}+1},$ (28)
@@
$\{P_{\mathcal{I},\mathcal{I}+1},P_{\mathcal{I},\mathcal{I}+2}\}\succeq-\eta(P_{\mathcal{I},\mathcal{I}+1}+P_{\mathcal{I}+1,\mathcal{I}+2}),$ (29)
@@
$\eta_{l}=||(\mathbb{I}-G_{1,\ldots,2l})(\mathbb{I}-G_{l+1,\ldots,3l})-(\mathbb{I}-G_{1,\ldots,3l})||,$ (30)
and $G_{i,i+1,...,i+m-1}$ denotes the projector onto the ground space of the original Hamiltonian $h$ on total $m$ sites from indices $i$ to $i+m-1$. Finally, using the above inequality, the bound on the gap of the original Hamiltonian is given as,
$\Delta_{M}(n=3l)=\frac{\epsilon_{2l}}{2}(1-2\eta_{l}),$ (31)
where $n$ denotes the number of sites in the original Hamiltonian. The lower bound of $\tilde{H}$ is $(1-2\eta)$ which is then rescaled by $\epsilon_{2l}/2$ to recover the bound for the original Hamiltonian $H$. Therefore, to prove the existence of a gap it is sufficient to find a value for $l$ that satisfies the condition $\eta_{l}<0.5$, which is useful in cases where analytical proofs are required [42, 32]. Given the form for the ground state projectors, it is possible to compute $\eta_{l}$ numerically by increasing the value of $l$ until the threshold of $\eta_{l}<0.5$ is achieved. The other possibility is to analytically bound $\eta_{l}$ as $\eta_{l}\leq\eta^{\prime}(l)$, where the function $\eta^{\prime}(l)$ becomes small enough for large $l$. This is problematic if the entire gap estimate (i.e. $\Delta_{M}$) is required, because, if the value of $l$ for which $\eta^{\prime}(l)<0.5$ is large, then it becomes harder to compute the prefactor $\epsilon_{2l}$ in eq. 31, as it requires the exact diagonalization of a $2l$-sites system.
III.2 Proof of inclusion

### 31 Mar 2026 v2

@@ 1 Introduction
Proving the existence of a spectral gap is a highly non-trivial mathematical challenge. Even for reasonable local Hamiltonians the task can be undecidable [19, 11, 18, 49]. Several techniques to lower bound the spectral gap have been developed for the better-behaved class of frustration-free Hamiltonians, i.e. systems in which the ground state minimizes each local term individually. The most notable lower bounds are based on the “martingale” method [44] and finite size criteria [32, 20]. More recently, methods inspired by the numerical bootstrap approach to conformal field theory have been proposed to determine spectral properties of simple quantum Hamiltonians as well as quantum many body systems [12, 13, 45].
@@
To put our results into context, we find it worthwhile to pause and briefly discuss the basic idea behind our method—and how it generalizes existing gap certification methods—on a slightly more technical level. The well known idea underlying all gap certification approaches is the fact that a frustration-free Hamiltonian $H$ has a gap greater or equal to $\delta$ if the operator $H^{2}-\delta H$ is positive semidefinite (i.e. all its eigenvalues are non negative). Existing methods thus all amount to finding ways to decompose $H^{2}-\delta H$ as a sum of positive terms. Usually a decomposition is chosen such that some terms are manifestly positive, while for others positivity can be proven given some positivity condition holds for an appropriately chosen local Hamiltonian—a finite size criterion. There are of course many possible choices of such decompositions, and finding the appropriate finite size criterion for a given system has thus far been an intricate task. The main idea of our method is that the search for a positive decomposition of $H^{2}-\delta H$ can be automated and solved in an algorithmic way. Namely we formulate the gap estimation problem as an optimization problem where $\delta$ is maximized under the constraint that $H^{2}-\delta H$ can be written as a sum of positive terms. A basic version of this idea has been proposed in Ref. [17]. Here we exhaust the search space in a systematic way through a hierarchy of optimization problems: at the $n$-th level of the hierarchy we optimize over all possible decompositions of $H^{2}-\delta H$ as a sum of translation-invariant $n$-body geometrically local terms.
The rest of the paper is organized as follows: In Section 2, we discuss the main technical ingredients of our method. In Section 3, we show how previous finite-size criteria arise as special cases of the semidefinite program introduced in Section 2. In Section 4, we present numerical results comparing lower bounds obtained using our method against those obtained with existing finite-size criteria.
2 The method
@@
Without loss of generality, we will shift the ground state energy to be 0, which, combined with frustration-freeness, implies both $H\succeq 0$ and $h_{i}\succeq 0$, where $\succeq$ is used to denote that the matrix is positive semidefinite.
The spectral gap is then the lowest non-zero eigenvalue of $H$ and we denote it by $\Delta_{N}$. Of particular relevance is the behavior of $\Delta_{N}$ as the system size increases. A model is defined to be gapped if the spectral gap does not close in the thermodynamic limit, that is $\Delta_{N\to\infty}=\Delta>0$.
@@
The above condition enforces that $H$ has no eigenvalues in the interval $(0,\delta)$. Let us stress that all the most commonly-used methods to prove the gap of a frustration free system rely on establishing condition (2) for a given value of $\delta$ [32, 7, 44, 31, 38, 39, 50].
The condition in eq. 2 allows us to formulate the gap estimation problem as an optimization problem, which turns out to be a semidefinite program (SDP).
$\Delta_{N}$ $=\max\delta\,\,$
@@
under the condition that the ground state energy of $H$ is zero. The optimal value of this SDP is exactly the energy of the first excited state of $H$. Additionally, at the optimal solution, both the ground state and the first excited state space of $H$ lie in the ground state space of $H^{2}-\Delta_{N}H$.
In the general case, solving this SDP is as hard as exactly diagonalizing the Hamiltonian, since this problem imposes a positivity constraint on a matrix of size $(d^{N}\times d^{N})$. As $N$ increases, very quickly it becomes computationally intractable to solve. In the upcoming section, we show how to relax the problem of computing exact spectral gaps. The resulting relaxation is a local, $n$-body, problem the solution to which is a lower bound on the gap of $N$ site periodic system for all $N\geq 2n$, and therefore holds for the infinite system as well.
2.1 Simplifications of the general SDP
@@
Geometric Locality condition. The first problem we need to address is that $H^{2}$ is not local: it consists of products of interactions terms $h_{i},h_{j}$ arbitrarily far apart. We therefore introduce a truncated operator denoted by $\left[H^{2}\right]_{n}$ which is $n$-local, by removing terms that act on spins farther than $n$ sites apart, namely
$\begin{split}\left[H^{2}\right]_{n}&=\sum_{|i-j|<n-k}h_{i}h_{j}\\ &=\sum_{i}\left(h_{i}^{2}+\sum_{i<j<i+n-k}\{h_{i},h_{j}\}\right)\end{split}$ (4)
(recall that $k$ is the locality range of the interaction term). If $n$ is chosen suitably (i.e. $n\geq 2k+1$), terms $h_{i}$ and $h_{j}$ with $j\geq i+n$ act on disjoint sets of spins and thus every discarded term is positive semidefinite (as a tensor product of positive terms). It follows that, if we define the following operator,
@@
where the the $i$ subscript indicates the position of the first spin on which $q_{n}$ acts (the $n$ subscript simply denotes the $n$-locality, i.e. that $q_{n}$ is supported on $n$ neighboring sites). If for some value of $\delta$, $Q_{n}(\delta)$ is positive semidefinite, we can hope to find such a positive generating term. It could, however, be the case that a local positive generating term $q_{n}$ can only be found for a lower value of $\delta$, and thus we are further relaxing the problem in this step. (In practice our numerical demonstrations below show that the effect of this relaxation step (eq. 6) is negligible.)
@@
We will now proceed to show how the condition (6) can be expressed in terms of local operators, and thus arrive at a finite size relaxation of section 2 that is numerically tractable for any finite support $n$.
2.2 Primal SDP
After substituting the operator $H^{2}$ in the exact SDP section 2 with $H^{2}]_{n}$, and imposing the condition that $Q_{n}(\delta)$ has a positive operator decomposition we arrive at the following problem
@@ 2.2 Primal SDP
where $Y_{n-1}$ can be any operator supported on n1n-1 sites. Condition (10) is chosen such that, when summed over $i$, the terms involving $Y_{n-1}$ cancel telescopically. It follows that the constraint (7) corresponds to looking for a suitable $Y_{n-1}$ that makes the local generating term $q_{n,i}$ positive semidefinite. In this way, we arrive at the final form of our SDP, which reads
@@
We will refer to the problem we arrived at in eq. 11 as the locally translationally invariant (LTI) gap SDP. It is a semidefinite program which involves a positivity constraint on an $n$-body (i.e. $d^{n}\times d^{n}$ dimensional) matrix. To obtain a positive decomposition of $Q_{n}(\delta)$, one option is to formulate it as a sum-of-squares (SOS) problem [36, 48, 47, 10, 46], where positivity is enforced through moment matrices associated with a chosen set of monomials. In contrast, we directly optimize the matrix $q_{n}$ itself, making use of an exact characterization of the degrees of freedom that leave the sum $Q_{n}(\delta)$ unchanged.
@@
2.3 Dual SDP
Any semidefinite program which is defined as a maximization has a dual formulation in terms of a minimization problem [14]. Importantly, the optimum of the dual serves as an upper bound on the primal optimal. When the two optima coincide, the problem exhibits so-called strong duality and one can equivalently solve either one of the problems. In this subsection we present the duals of the spectral gap SDPs introduced in the previous subsections and discuss several useful insights which become transparent when the problem is viewed from the perspective of the dual.
The dual program of the exact spectral gap SDP of section 2 is
@@ 2.3 Dual SDP
In view of eq. 15, we can now justify the first relaxation step eq. 4, in which we truncated $H^{2}$ by omitting the terms $h_{i}\otimes h_{j}$ for |ij|>n|i-j|>n. We argue that the contribution of such highly nonlocal terms to the objective in eq. 15 can be, in some cases, negligible. This is because of the local nature of excitations in gapped systems. As shown in Ref. [26], an excited state $\psi_{p,\alpha}$} with momentum $p$ and energy $E_{\alpha}$, which is separated by an energy gap from above and from below, can be well approximated by a momentum superposition of an $l$-local operator $O^{(l)}$ acting on the ground state $\psi_{0}$:
[…]
The dual of the LTI gap SDP eq. 11 has the following form
$\delta_{\mathrm{LTI}}(n)=\min_{\rho_{n}}$ Tr[$\;\Tr\left[\rho_{n}\left(h_{1}^{2}+\sum_{j=2}^{n-1}\{h_{1},h_{j}\}\right)\right]$ (17)
@@
where the optimization variable $\rho_{n}$ is the reduced state on $n$ consecutive sites. The constraint
@@
indeed requires $\rho_{n}$ to obey local translation invariance and it arises as a dual to the constraint eq. 12.
Due to strong duality, the following relation is satisfied between the primal and dual optimal variables in eqs. 11 and 17 respectively
@@
This relation which is known as complementarity slackness has several useful consequences. Among others, it provides us with the gradient of the solution to the SDP with respect the parameters defining the problem data, without incurring further computational cost [52]. In our case the problem data is the Hamiltonian interaction term $h$ and, similarly to the Hellmann–Feynman theorem, the gradient with respect to $h$ of the solution to eq. 17 can be computed from the minimizing state:
@@
This could be used to maximize the gap over families of parent Hamiltonians of a given ground state. We elaborate on this in Appendix C.
3 Connection to previous finite-size criteria
In this section, we compare our method to finite-size criteria-based techniques, which are the prevalent methods used to lower bound spectral gaps. We find that the LTI SDP eq. 11 will always detect a gap greater or equal to the ones detected by existing finite size methods. This is simply because each finite size method looks for a specific $n$-local positive decomposition, whereas our SDP optimizes over all such decompositions. In this section, we prove that the decompositions corresponding to different finite size methods are feasible (yet not necessarily optimal) in our SDP. In other words, they are included in our search. Additionally, the finite-size method requires the interaction terms in the Hamiltonian to be projectors, whereas the LTI SDP works directly with the original interactions. While using the projector form does not affect the condition for the existence of the gap, it reduces the accuracy of the resulting bound.
Note that the lower bounds based on finite size criteria pertain to the spectral gaps $\Delta_{N}$ of systems with periodic boundary. Their expressions, however, involve the computation of spectral gaps for open boundary Hamiltonians, which we denote as $\epsilon_{n}$. For consistency with the formulation of the existing methods [32, 24], we restrict the discussion to nearest-neighbor projector Hamiltonians
@@ 3 Connection to previous finite-size criteria
Importantly, if the local term $h$ of a given frustration-free Hamiltonian is not a projector, it can always be modified into a projector Hamiltonian by adjusting the eigenvalues of the subspace orthogonal to the ground space of $h$. It follows from a simple argument that proving the gap of this second Hamiltonian implies a gap for the first as well [24](c.f. Section 4). We will now prove that the set of feasible local terms in the LTI-SDP eq. 11 includes the ones obtained from known finite size criteria. Thus, the value of LTI SDP achieves a tighter lower bound on the spectral gap.
3.1 Finite-size criteria
@@
(Method 1) Based on finite system gaps. First formulated by Knabe [32], these criteria provide a way to relate the spectral gap of a finite open-boundary system to the spectral gap in the thermodynamic limit. The open boundary Hamiltonian on a finite system of size $n$ is,
$H_{n}^{\mathrm{OBC}}=\sum\limits_{i=1}^{n-1}h_{i},$ (21)
If $H_{n}^{\mathrm{OBC}}$ has a spectral gap of $\epsilon_{n}$  [^1] Note that for ease of reading, we denote spectral gaps of periodic boundary Hamiltonians with $\Delta$ and of open boundary Hamiltonians with $\epsilon$., then the lower bound to the gap in the thermodynamic limit is given as
$\delta(n)=c_{n}\left(\epsilon_{n}-t_{n}\right).$ (22)
Here $c_{n}$ is a function of $n$ which does not affect the condition for existence of the gap. The quantity $t_{n}$ defines a threshold, i.e., if the local gap satisfies the condition $\epsilon_{n}>t_{n}$, then the system is gapped in the thermodynamic limit. If $\Delta_{m}$ is the gap of periodic system on $m$ sites, then the Knabe bound [32] is
$\Delta_{m}\geq\delta_{\mathrm{Knabe}}(n)=\frac{n-1}{n-2}\left(\epsilon_{n}-\frac{1}{n-1}\right)$ (23)
@@
$\Delta_{m}\geq\delta_{\mathrm{GM}}(n)=\frac{5}{6}\frac{n^{2}+n}{n^{2}-4}\left(\epsilon_{n}-\frac{6}{n^{2}+n}\right)$ (24)
for $m>2n>4$. Comparing the bounds with eq. 22, the functions $c_{n}$ and $t_{n}$ satisfy the following properties
$\lim_{n\to\infty}c_{n}\leq 1,\quad\lim_{n\to\infty}t_{n}=0.$ (25)
@@
This method was introduced by Nachtergaele, Fannes, and Werner in Ref. [44, 20], and is also known as the “martingale” criterion. Instead of using the gap of a finite size system, the presence of a gap is established by checking if a relation in terms of the ground space projectors is satisfied. The bound is obtained by first constructing a new coarse-grained Hamiltonian $\tilde{H}=\sum_{\mathcal{I}}\tilde{h}_{\mathcal{I},\mathcal{I}+1}$ obtained by blocking $l$ sites into a single site, i.e.
$\mathcal{I}=\{(i-1)l+1,...,il\}$ (27)
$\tilde{h}_{\mathcal{I},\mathcal{I}+1}=\sum\limits_{j=(\mathcal{I}-1)l+1}^{(\mathcal{I}+1)l-1}h_{j,j+1}.$
The locality of the new Hamiltonian $\tilde{h}_{\mathcal{I},\mathcal{I}+1}$ is $2l$-local in the original system. Using eq. 27, we have the following relation between the original and the blocked Hamiltonian, as in the latter every term form the former appears either once or twice:
$H\succeq\frac{\tilde{H}}{2}.$ (28)
W.l.o.g, the coarse-grained Hamiltonian $\tilde{h}$ can be converted into the corresponding projector Hamiltonian $K=\sum_{\mathcal{I}}P_{\mathcal{I},\mathcal{I}+1}$, which satisfies the relation,
$\tilde{h}_{\mathcal{I},\mathcal{I}+1}\succeq\epsilon_{2l}P_{\mathcal{I},\mathcal{I}+1},$ (29)
@@
$\{P_{\mathcal{I},\mathcal{I}+1},P_{\mathcal{I}+1,\mathcal{I}+2}\}\succeq-\eta_{l}(P_{\mathcal{I},\mathcal{I}+1}+P_{\mathcal{I}+1,\mathcal{I}+2}),$ (30)
@@
[…]
and $G_{i,i+1,...,i+m-1}$ denotes the projector onto the ground space of the $m$-site Hamiltonian with open boundary conditions
$H_{m,i}^{\mathrm{OBC}}=\sum_{j=i}^{i+m-1}h_{j}.$ (32)
Finally, using the above inequality, the bound on the gap of the original Hamiltonian is given as,
$\delta_{\mathrm{M}}(3l)=\frac{\epsilon_{2l}}{2}(1-2\eta_{l}).$ (33)
The lower bound on gap of $K$ is $(1-2\eta_{l})$ which is then rescaled by $\epsilon_{2l}/2$ to recover the bound for the original Hamiltonian $H$. Therefore, to prove the existence of a gap it is sufficient to find a value for $l$ that satisfies the condition $\eta_{l}<0.5$, which is useful in cases where analytical proofs are required [15, 27]. Given the form for the ground state projectors, it is possible to compute $\eta_{l}$ numerically by increasing the value of $l$ until the threshold of $\eta_{l}<0.5$ is achieved. The other possibility is to analytically bound $\eta_{l}$ as $\eta_{l}\leq\eta^{\prime}(l)$, where the function $\eta^{\prime}(l)$ becomes small enough for large $l$. This is problematic if the entire gap estimate (i.e. $\delta_{\mathrm{M}}$) is required, because if $\eta^{\prime}(l)<0.5$ is only reached for a large value of $l$, then it becomes harder to compute the prefactor $\epsilon_{2l}$ in eq. 33, as it requires the exact diagonalization of a $2l$-sites system.
3.2 Proof of inclusion